OpenStudy (anonymous):

what is the product of 3sqrt _2b 49sqrt-b)

OpenStudy (anonymous):

\[\sqrt{3}(2b \sqrt{49}-b)\]

OpenStudy (anonymous):

b sqrt3 times 7 sqrt 3

OpenStudy (anonymous):

\[7b \sqrt{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thnx